k2(m-a) divided by x = x
solve "a"

Answer by stanbon(75887)

k2(m-a) divided by x = x
solve "a"
------------
[k^2(m-a)]/x = x
------
[k^2(m-a)] = x^2
---
m-a = x^2/k^2
---
a = m - [x/k]^2
-------------
Cheers,
Stan H.
----------------